# Large Language Models as Automatic Annotators and Annotation Adjudicators for Fine-Grained Opinion Analysis

### v1

@@ 4.1 Problem Statements
LLM as opinion annotators Given the input text $Ti$ from corpora $\mathcal{T}$, our objective is to extract all expressed opinions $Oi=\{o_{1},o_{2},..o_{n}\}$. The elements of single opinion ($o_{j}$) is task dependant:
  • ASTE opinions are formulated as the triple $o_{j}=\{at,s,op\}$, where $at$ is the aspect term, $s$ is the sentiment and $op$ is the opinion term.
  • ACOS opinions are quadruples $o_{j}=\{at,s,op,ac\}$, where $at$, $s$ and $op$ encode similar concepts as ASTE and $ac$ is aspect category made up of entity $E$ and attribute $A$ (expressed as E#AE\#A).
@@ 6 Results
The lower overall scores highlight the greater challenge of annotating quadruples, a task further complicated by domain differences: the laptop domain, with 114 distinct categories $ac$, proved more difficult than the restaurant domain, with only 12, adversely affecting the exact-match F1 score on the full quadruples.
@@ 7 Analysis
In Table 4, sentiment polarity ($s$) predictions have the highest alignment with respect to human annotations. There are fewer $at$ and $op$ spans than $s$, but they are very comparable in alignment. The AT extraction and assignment of correct polarity (s&ats\&at) is more challenging than individually labelling $s$ and $at$, but it is less challenging than $s$ and $op$ which is expected as sentiment polarity and words used to express the polarity are easier to predict jointly. Based on these scores, the most challenging step short of predicting the entire triplet jointly is at&opat\&op identification (i.e. correctly identifying which $op$ span expresses the $op$ on which target.
Similar trends to the ASTE triplets can be found in Table 5. All the models short of 32B exhibit high domain disparity, which can be clearly attributed to the previously discussed difficulty of correctly predicting $ac$.
@@ 7.2 Error Analysis
In the Table 7, we observe how the entire prediction in ASTE is deemed incorrect in triplet evaluation due to a minor mistake in identifying the aspect term ($ac$). In the ACOS formulation for the same example, we see the challenge model faces in identifying $at$.
@@
I have eaten here three times and have found the quality and variety of the fish to be excellent. $s$: Positive $at$: quality $op$: excellent $s$: Positive $at$: fish $op$: excellent $s$: Positive $at$: quality $ac$: food#quality $op$: excellent $s$: Positive $at$: variety $ac$: food#style $op$: excellent $s$: Positive $at$: fish $ac$: food#quality $op$: excellent $s$: Positive $at$: fish $ac$: food#style $op$: excellent
All their menu items are a hit, and they serve mimosas. $s$: Positive $at$: menu items $op$: hit $s$: Positive $at$: menu items $op$: hit $s$: Positive $at$: menu items $ac$: food#quality $op$: hit $s$: Neutral $at$: mimosas $ac$: drinks#general $op$: null $s$: Positive $at$: menu items $ac$: food#quality $op$: hit $s$: Positive $at$: null $ac$: drinks#style $op$: hit

### v2

@@ 4.1 Problem Statements
LLM as opinion annotators Given the input text TiTi from corpora $\mathcal{T}$, our objective is to extract all expressed opinions $Oi=\{o_{1},o_{2},..o_{n}\}$. The elements of single opinion ($o_{j}$) is task dependant:
  • ASTE opinions are formulated as the triple $o_{j}=\{at,s,op\}$, where atat is the aspect term, $s$ is the sentiment and opop is the opinion term.
  • ACOS opinions are quadruples $o_{j}=\{at,s,op,ac\}$, where atat, $s$ and opop encode similar concepts as ASTE and acac is aspect category made up of entity $E$ and attribute AA (expressed as E#AE\#A).
@@ 6 Results
The lower overall scores highlight the greater challenge of annotating quadruples, a task further complicated by domain differences: the laptop domain, with 114 distinct categories acac, proved more difficult than the restaurant domain, with only 12, adversely affecting the exact-match F1 score on the full quadruples.
@@ 7 Analysis
In Table 4, sentiment polarity ($s$) predictions have the highest alignment with respect to human annotations. There are fewer atat and opop spans than $s$, but they are very comparable in alignment. The AT extraction and assignment of correct polarity (s&ats\&at) is more challenging than individually labelling $s$ and atat, but it is less challenging than $s$ and opop which is expected as sentiment polarity and words used to express the polarity are easier to predict jointly. Based on these scores, the most challenging step short of predicting the entire triplet jointly is at&opat\&op identification (i.e. correctly identifying which opop span expresses the opop on which target.
Similar trends to the ASTE triplets can be found in Table 5. All the models short of 32B exhibit high domain disparity, which can be clearly attributed to the previously discussed difficulty of correctly predicting acac.
@@ 7.2 Error Analysis
In the Table 7, we observe how the entire prediction in ASTE is deemed incorrect in triplet evaluation due to a minor mistake in identifying the aspect term (acac). In the ACOS formulation for the same example, we see the challenge model faces in identifying atat.
@@
I have eaten here three times and have found the quality and variety of the fish to be excellent. $s$: Positive atat: quality opop: excellent $s$: Positive atat: fish opop: excellent $s$: Positive atat: quality acac: food#quality opop: excellent $s$: Positive atat: variety acac: food#style opop: excellent $s$: Positive atat: fish acac: food#quality opop: excellent $s$: Positive atat: fish acac: food#style opop: excellent
All their menu items are a hit, and they serve mimosas. $s$: Positive atat: menu items opop: hit $s$: Positive atat: menu items opop: hit $s$: Positive atat: menu items acac: food#quality opop: hit $s$: Neutral atat: mimosas acac: drinks#general opop: null $s$: Positive atat: menu items acac: food#quality opop: hit $s$: Positive atat: null acac: drinks#style opop: hit